OpenStudy (anonymous):

Find an equation of the line having the given slope and containing the given point. Express your answer in the given form x=a,y=b or y=mx+b m=-6(4,0)

OpenStudy (anonymous):

0 = -6(4) + b 0 = -24 + b 24 = b y = -6x + 24